OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. x + y = 10 x – y = 2

OpenStudy (anonymous):

would this be 8+2=10 10-8=2

OpenStudy (anonymous):

add the two equations to each other...

OpenStudy (anonymous):

x+y=12 is this right

OpenStudy (anonymous):

x+x =? y-y =?

OpenStudy (anonymous):

8+10=18 2-8=-6

OpenStudy (anonymous):

x + y =10 + x - y = 2 ---------------

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

ask a question.

OpenStudy (anonymous):

i don't understand what to do after this

OpenStudy (anonymous):

add them. x+x = ? y-y =? 10+2 =?

OpenStudy (anonymous):

the last part is 12 but i dont know what to put in the x and y spot

OpenStudy (anonymous):

does it all have to equal 12

OpenStudy (anonymous):

what if x was 4? what would x+x equal?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

yes what if x was 3? what would x+x equal?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

see the idea? we don;t have to know what x is.. adding x to x is 2x no matter whether x is 3 or 4 or 1000

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

ok thanks so the answer is 2x and -2y

OpenStudy (anonymous):

2x part is right how about y -y?

OpenStudy (anonymous):

would there be a y at all

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so 2x=12 is all that is left x=?

OpenStudy (anonymous):

so the answer is 6

OpenStudy (anonymous):

x=6 ... now use that value to go back and find y 6+y =10 y=?

OpenStudy (anonymous):

so y would equal 4

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thank you so much i love this website

OpenStudy (anonymous):

so it would be no solutions right

OpenStudy (anonymous):

because 6 and 4 are not equal

OpenStudy (anonymous):

x=6 and y=4 is the answer